A mystery number is subtracted from 500. The answer has three consecutive digits and the digits of the answer add to 12. What is the mystery number?
Found 2 solutions by ikleyn, MathTherapy:
Answer by ikleyn(52800)

.
There are more than one different solutions/answers.


One, the most simple solution is  500 - x = 345.    ====>  x = 500 - 345.


The other is 500 - x = 543  ====>  x = 500 - 543 = - 43. 


-------------


There is nothing in the condition (written quite unprofessionally) that prohibit to have many other solutions.


Notice that "consecutive digits" does not mean "consecutive numbers". It has totally different meaning.


Answer by MathTherapy(10552)

A mystery number is subtracted from 500. The answer has three consecutive digits and the digits of the answer add to 12. What is the mystery number?
The lady's correct.
Unless other SPECIFIC clues are given, the result, after the subtraction from 500, could be one of 6 different numbers: 345, 543, 534, 453, 435, or 354.
This makes the "mystery" number: